OpenStudy (anonymous):

olve: 5(x + 6) less than or greater to 3x - 4 x greater than or equal to -17 x greater than or equal to 17 x less than or greater to -17 x less than or greater to -5

OpenStudy (highschoolmom2010):

5x+30=3x-4 -30 -30 5x=3x-34 -3x -3x 2x=-34 /2 /2 x=-17

OpenStudy (anonymous):

greater or less ?